When a sample of baking soda is strongly ignited in a crucible ,it suffered a loss in weight of 3.1g the mass of baking soda is?

Grade:12th pass

5 Answers

Vikas TU
7 years ago
The reacn. is: 
   2NaHCO3     ---------->            Na2CO3     +      CO2        +    H2O
     => 2* 84 = 168 gm.
 
CO2 and H2O => 44 + 18 => 62 gm.
Just divide them we get,
=> 168/62 = 8 gm approx.
 
As 3.1 g was Na2CO3 gm.                   
Shravani
7 years ago
If we divide 168/62 the answer is not 8 its like 2 (approx) and can u answer me the question by seeing again
Vikas TU
7 years ago
Correction:
mass of baking soda => 168 x 3.1 / 62 => 8 gms approx.
mutiply 3.1 gm with baking soda mass too.
Debanko Kundu
6 years ago
2NaHCO3 ------>NaCO3 + H2O + CO2mass of NaHCO3 = 2 × 84 =168mass of CO2 +H2O = 62Mass lost = 3.1 Mass of NaHCO3 = 168 × 3.1 / 62 = 8.4
Satya boppana
4 years ago
168 grams = weight of 2 NaHCo3
It losses 62 grams on decomposition by producing water and co2
168=62
?=3.1
Ans: 31×84÷31×10
8.4 grams